Roger sees water balloons fall past his window. He notices that each balloon strikes the sidewalk 0.59 s after passing his window. Roger's room is on the second floor, 10 m above the sidewalk.


How fast are the balloons traveling when they pass Roger's window?



Assuming the balloons are being released from rest, from what floor are they being released? Each floor of the dorm is 5.0 m high.


1
Expert's answer
2022-03-01T17:33:15-0500

"v_2=v+gt,"

"v_2=\frac{2h}g-v,"

"v=\frac hg-\frac{gt}2=1.87~\frac ms,"

"v_2=7.7~\frac ms,"

"H=\frac{v_2^2}{2g}=30~m,"

"n=\frac H5=6" floor.